A rod of mass m and length 2R is fixed along the diameter of the ring of same mass m and radius R. the combined body is rolling without slipping along x-axis. Then the angular momentum about z-axis is having magnitude equal to  

Grade:12th pass

1 Answers

Fahad
6 years ago
For rod, L1= sum of orbital angular momentum Lo of cm with respect to origin or centre and spin angular momentum Lcm=> Lo + Lcm= mv*R + I.w= mvr + [mv (2R)^2 / 12] . was there`s no slipping then w= v/R= mvr + mR^2/3 .(v/R)= mvr + mvr/3 = 4mvr/3For ring total angular momentum=2mvr Therefore total angular momentum of system is, => 2mvr + 4mvr/3= 10mvr/3